Session: Mar 2007 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (TaoAn at Tue Mar 6 14:25:39 2007) Success :

Medicina (orfei at Tue Mar 6 15:05:14 2007) Success :

Metsahovi (Ari Mujunen at Tue Mar 6 16:02:46 2007) Success :
Weather deteriorated early in the morning to be very foggy and humid; surprisingly that is seen in Tsys figures much later, at around UT 9.

Onsala (Michael Lindqvist at Tue Mar 6 16:34:50 2007) Minor failures :
Bad weather, heavy rain. Clock problem: 2007.065.12:31:22.41?ERROR ma -4 Device fm timed-out on response. 2007.065.12:31:28.36?ERROR ma -8(fm) 2007.065.12:31:28.68?ERROR sc -2 setcl: formatter time out-of-range 2007.065.12:31:28.93?ERROR sc -2 setcl: formatter time out-of-range 2007.065.12:31:28.93?ERROR sc -10 setcl: failed too many times, couldn't check formatter time 2007.065.12:31:28.93?ERROR sc -12 setcl: FS to computer time difference 0.5 seconds or greater 2007.065.12:31:28.93?ERROR sc -15 setcl: Cannot determine NTP sync state

Effelsberg (A. Polatidis at Thu Mar 8 14:24:46 2007) Success :
Telescope slewing and settling on source resulted in 10sec or less integration ON source. No other problems

Noto (P. Cassaro at Sun Mar 11 01:59:04 2007) Minor failures : ANT
Possible poor gain due to problems with active surface and sub-reflector.

Jodrell_Bank (Alastair Gunn at Tue Mar 20 14:42:45 2007) Minor failures : WIND
The MkII telescope was used (Jb2). Telescope parked due to high winds from start of experiment until 03:35 UT. Missing PRC file at start, corrected within five minute. No other known problems.